# CaveAgent: Transforming LLMs into Stateful Runtime Operators

### v1

@@ 5 Conclusion
We present CaveAgent, a novel framework that transforms LLM tool use from stateless JSON function calling to persistent, object-oriented stateful runtime management. CaveAgent enables agents to maintain high-fidelity memory of complex objects and execute sophisticated logic via Python code. Experiments on Tau2-bench show that this approach significantly outperforms SOTA baselines in multi-turn success rates (+10.5%) and token efficiency. Crucially, the programmatic verifiability of the runtime state provides a rigorous ground for future advancements in Reinforcement Learning and runtime-mediated multi-agent coordination, marking a critical step towards more reliable and capable autonomous agents. Qualitative case studies are provided in Appendix F.
6 Related Work
6.1 Tool Learning & Instruction Following (JSON-centric Paradigm)
The foundational approach to equipping Large Language Models (LLMs) with agency has relied on a "Classification-Slot Filling” paradigm, where models interface with external environments via structured data formats, predominantly JSON. Seminal works such as ToolLLM (qin2023toolllm) and Gorilla (patil2024gorilla) demonstrated that LLMs could be fine-tuned to navigate massive API indices and mitigate hallucinations by strictly adhering to predefined schemas. This structured interaction model was further formalized by industry standards like GPT-4 Function Calling and adopted by agentic frameworks such as ReAct (yao2022react) and the JSON mode of AutoGen (wu2023autogen), which orchestrate reasoning through iterative schema population. Nevertheless, the JSON-centric paradigm imposes severe architectural constraints. First, JSON is inherently a static data interchange format lacking native control flow; it cannot represent loops or conditional logic, forcing agents into expensive, multi-turn interactions to execute complex workflows (wang2024executable). Second, the verbose syntax of JSON introduces significant token overhead, resulting in low information density and high latency (wang2024executable). Finally, the rigidity of schema enforcement creates a fragility trade-off, where complex nested structures increase the probability of syntax errors and hallucination (patil2024gorilla).
@@ 6.2 Code as Action & Programmatic Reasoning
Recognizing the limitations of static schemas, recent research has explored "Code as Action” paradigm, where executable code (primarily Python) serves as the unified medium for reasoning and tool invocation. wang2024executable challenged the JSON-Schema convention with CodeAct, proposing executable Python code as a unified representation for both reasoning and action. Their work demonstrated that code-based interactions reduce multi-turn overhead by up to 30% compared to JSON-based methods and improve task success rates by 20% on benchmarks like M3ToolEval. This shift utilizes the Turing-complete nature of code to naturally express complex logic, loops, and variable dependencies that are cumbersome in JSON. This paradigm extends to domain-specific reasoning. suris2023vipergpt introduced ViperGPT, which composes vision modules into executable subroutines to solve visual queries, rendering the reasoning process interpretable. Similarly, chen2022program and gao2023pal proposed Program of Thoughts (PoT) and Program-aided Language Models (PAL), respectively. These frameworks decouple computation from reasoning by delegating arithmetic and symbolic logic to a Python interpreter, thereby mitigating the calculation errors common in pure language models. The efficacy of these methods is further amplified by code-optimized models such as DeepSeek-Coder-V2 (zhu2024deepseek), which exhibit superior performance in following complex programmatic instructions.
@@ 6.3 Context Management & Stateful Architectures
The constraints of LLM context windows have necessitated advanced memory management strategies. packer2023memgpt introduced MemGPT, which implements an OS-inspired virtual context management system, organizing memory into tiers (main vs. external) to handle long-horizon tasks. Similarly, qiao2023taskweaver proposed TaskWeaver, a code-first framework that attempts to maintain stateful execution for data analytics by preserving data structures like DataFrames across turns. However, existing approaches largely rely on Retrieval-Augmented Generation (RAG) or textual summarization to manage context. These methods are inherently lossy: converting complex runtime objects (e.g., high-dimensional matrices, class instances) into text or vector embeddings strips them of their structural integrity and executable properties. CaveAgent addresses this by proposing Runtime-based Context Compression. Unlike prior work that externalizes state to vector stores, we utilize Variable Injection to treat the Python runtime itself as a high-fidelity external memory. This allows arbitrary variables to be persisted in their native object form, maintaining full manipulability without the overhead of re-tokenization or the information loss associated with serialization.
6.4 Multi-Agent Coordination Mechanisms
Research into multi-agent systems has focused on structuring collaboration through natural language communication. li2023camel proposed CAMEL, a role-playing framework that facilitates autonomous cooperation via communicative agents. Building on this, qian2023chatdev introduced ChatDev, which organizes agents into a "chat chain" following a waterfall software development model, while hong2023metagpt developed MetaGPT, which encodes Standardized Operating Procedures (SOPs) into agent prompts to streamline complex workflows. However, these frameworks predominantly rely on text-based message passing for coordination. This architecture introduces a critical serialization bottleneck: transferring complex state (e.g., a trained machine learning model or a processed dataset) between agents requires converting it into natural language descriptions or intermediate files, leading to high latency and potential ambiguity. CaveAgent overcomes this via Runtime-Mediated State Flow. By leveraging the shared variable space established in our runtime architecture, agents collaborate by directly injecting and retrieving variables. This shifts the coordination paradigm from "communication by talking" to "communication by shared state," enabling atomic, lossless, and zero-latency information exchange.
@@ 7 Acknowledgment
We thank Rui Zhou, a professional UI designer affiliated with Hong Kong Generative AI Research & Development Center, HKUST, for his professional contributions to the figure design of this paper.
@@ Appendix B What Happens in Semantic Stream
The following sections detail the exact prompt templates used to instruct the Semantic Stream in CaveAgent to help readers understand what happens in this stream. The system prompt is dynamically constructed by combining the Agent Identity, Context Information (functions, variables, types), and specific Instructions.
@@ Appendix C What Happened in Runtime Stream
While the Semantic Stream governs the agent’s reasoning and planning, the Runtime Stream serves as the system’s execution engine and persistent memory. This stream operates as a dedicated Python kernel where the actual “work” of the agent (data manipulation, tool invocation, and state transitions) occurs. The interaction between the two streams follows a strict chronological topology, synchronized through an interleaved exchange of code instructions and execution feedback.
@@ C.1 Environment Initialization via Injection
The runtime lifecycle begins not with an empty state, but with Context Injection. Before the reasoning cycle commences, the user (or the system orchestration layer) initializes the runtime environment by injecting native Python objects directly into the global namespace.
@@
  • Variable Injection: Domain-specific data—such as complex DataFrames, graph structures, or class instances—are instantiated within the “memory" in the runtime stream.
@@ C.3 Illustrative Case Study
The analogy to view the runtime-stream as a jupyter-notebook with multiple cells (where each cell corresponds to the execution of Runtime Stream of each time step) could help us understand the mechanism of stateful management, especially how the states remain persistent across each cell.
@@ D.1 Type Proficiency Cases
The Type Proficiency category evaluates the agent’s competency in precise, state-aware manipulation of Python runtime elements. Unlike generic code generation, this section rigorously tests the agent’s “working memory” across three structural tiers: Simple Types (primitives types such as list, dictionary and string), Object Types (custom classes), and Scientific Types (high-dimensional complex data). Mastery of these domains serves as the foundational prerequisite for complex reasoning tasks.
@@ D.3 Multi-turn Cases
This class of test cases is designed to evaluate the agent’s capability to process complex, sequential instructions and maintain state precision over long-horizon scenarios. Unlike single-turn tasks where information is self-contained, these scenarios require the agent to maintain a persistent memory of the system’s status, as subsequent queries often depend on the outcome of previous actions. We categorize these multi-turn benchmarks into two distinct domains: Smart Home Control and Financial Account Management.
@@ D.3.1 Smart Home
As illustrated in Figure 11, the weekend_party case spans a simulated 24-hour cycle. The agent must maintain a coherent environment state—transitioning from a quiet morning to a loud party and finally to a secure night mode, without drifting from the user’s cumulative intent.
@@ D.3.2 Financial Account
The Financial Account benchmark evaluates the agent’s capability to maintain strict numerical integrity and execute state-dependent logic within a banking ledger system. Unlike the relative adjustments in Smart Home, this domain demands exact integer arithmetic, where the agent must process a continuous stream of transactions—including deposits, interest applications, and loan amortizations—without cumulative drift.
This scenario imposes two critical constraints designed to stress-test the agent’s reasoning stability:
@@ Appendix E Stateful Runtime-Mediated Multi-Agent Coordination
The function calling paradigm in CaveAgent introduces three foundational innovations for multi-agent coordination; Figure 1 illustrates an intuitive and straightforward example of these implications. In this paper, we primarily focus on qualitative analysis and provide intuitive case studies to facilitate understanding, leaving rigorous methodological development and quantitative justification for future work. We introduce the high-level idea below.
@@ E.1 Meta-Agent Runtime Control
Sub-agents are injected as first-class objects into an Meta-agent’s runtime, enabling the Meta-agent to programmatically access and manipulate child agent states through generated code. Rather than following predefined communication protocols, the Meta-agent dynamically sets variables in sub-agent runtimes, triggers execution, and retrieves results, enabling adaptive pipeline construction, iterative refinement loops, and conditional branching based on intermediate states.
@@ E.2 State-Mediated Communication
Inter-agent data transfer bypasses message-passing entirely. Agents communicate through direct runtime variable injection: the Meta-agent retrieves objects from one agent’s runtime and injects them into another’s as native Python artifacts (DataFrames, trained models, statistical analyses), preserving type fidelity and method interfaces without serialization loss.
@@ E.3 Shared-Runtime Synchronization
For peer-to-peer coordination, multiple agents can operate on a unified runtime instance, achieving implicit synchronization without explicit messaging. When one agent modifies a shared object, all peers perceive the change immediately through direct reference. New entities injected into the shared runtime become instantly discoverable, enabling emergent interaction and collaborative manipulation of a unified "world" model with low coordination overhead.
How town simulation demonstrates this capability When the Meta-agent modifies the weather state, all resident agents observe the change through direct attribute access; when a new location and manager are injected, existing agents can immediately query and interact with them.
Together, these patterns transform multi-agent systems from lossy text-based message exchange into typed, verifiable state flow, and enable automated validation of inter-agent handoffs and seamless integration with downstream pipelines.
Appendix F Features
F.1 Case Analysis in Tau2-bench
To empirically validate the architectural advantages of CaveAgent, we analyzed trajectory differences on the tau2-bench retail benchmark. CaveAgent achieved a 72.8% success rate (83/114) compared to 62.3% (71/114) for the baseline JSON agent (Kimi K2 backbone), yielding a 10.5% improvement. We conducted a root cause analysis on the 24 tasks where CaveAgent succeeded but the baseline failed.
F.1.1 Failure Taxonomy of the Baseline
@@
F.1.2 Architectural Advantages: Loops and Conditionals
The analysis reveals that CaveAgent’s superiority stems from its ability to generate programming constructs, specifically loops (used in 92% of winning cases) and conditionals (83%), which resolve the semantic gaps inherent in single-step function calling.
@@ F.1.2 Architectural Advantages: Loops and Conditionals
F.2 Smart Home
Figure 14 illustrates the mechanistic advantage of CaveAgent through a toy smart-home example. The architecture separates the Semantic Stream (logic generation) from the Runtime Stream (state storage). This design enables two critical capabilities absent in standard JSON agents:
@@ F.2 Smart Home
Figure 14: Demonstration of CaveAgent in a toy Smart-home example: the Semantic Stream interact with the Runtime stream via generating code, manipulating the stateful objectives (variables) in the persistent runtime.

### v2

@@ 4 Conclusion
We present CaveAgent, a framework that transforms LLM tool use from stateless JSON function calling to persistent, object-oriented stateful runtime management. CaveAgent enables agents to maintain high-fidelity memory of complex objects and execute sophisticated logic via Python code. Experiments on Tau2-bench show that this approach outperforms SOTA baselines in multi-turn success rates (+10.5%) and token efficiency. Beyond performance gains, a key contribution is the programmatic verifiability enabled by CaveAgent’s architecture: because runtime state is deterministically inspectable, agent behavior can be evaluated automatically without human annotation, establishing a structural foundation for Reinforcement Learning with Verifiable Rewards and runtime-mediated multi-agent coordination. Qualitative case studies are provided in Appendix G.
Limitations.
CaveAgent requires tools to be Python-callable or wrapped as Python functions; tools with non-Python interfaces (e.g., native binaries, proprietary REST APIs) require manual wrapper creation. The persistent runtime consumes memory proportional to the complexity of stored objects; for extremely long sessions with large data artifacts, memory management becomes a concern. While agent behavior is in principle programmatically verifiable, designing comprehensive fine-grained benchmarks for stateful evaluation remains future work. Finally, the multi-agent coordination capabilities are demonstrated qualitatively; rigorous quantitative evaluation of runtime-mediated multi-agent systems is left for future investigation.
5 Background and Related Work
5.1 Tool Learning & Function Calling
The foundational approach to LLM tool use relies on a JSON-centric paradigm. The ReAct framework (yao2022react) generates actions as free-text strings requiring heuristic parsing, making it prone to hallucination and format errors (patil2024gorilla, liu2023agentbench). To address this, JSON-Schema function calling (patil2024gorilla, qin2023toolllm) constrains action generation to structured schemas, formalized by GPT-4 Function Calling and adopted by frameworks such as AutoGen (wu2023autogen). Constrained decoding methods like xGrammar (dong2024xgrammar) can further enforce syntactically valid JSON at low overhead. However, format correctness does not resolve deeper architectural limitations: JSON is a static interchange format lacking native control flow, its verbose syntax incurs token overhead (wang2024executable), and rigid schemas increase the probability of hallucination on complex nested structures (patil2024gorilla). The entire pipeline operates as a text-based serialization loop: the schema is flattened into the prompt, the LLM predicts a JSON payload, and middleware serializes results back into text, inheriting context explosion and error propagation (packer2023memgpt, wang2023mint).
@@ 5.2 Code as Action & Programmatic Reasoning
Recognizing these limitations, the “Code as Action” paradigm uses executable Python as a unified medium for reasoning and tool invocation. wang2024executable proposed CodeAct, reducing multi-turn overhead by up to 30% and improving task success rates by 20%. This shift leverages the Turing-complete nature of code to express loops, conditionals, and variable dependencies. The paradigm extends to domain-specific reasoning: ViperGPT (suris2023vipergpt) composes vision modules into executable subroutines, while Program of Thoughts (chen2022program) and PAL (gao2023pal) delegate arithmetic and symbolic logic to a Python interpreter. However, current code agents retain architectural limitations: CodeAct provides no explicit APIs for external object injection or retrieval, relying on a textualization bottleneck where intermediate states must be serialized via print and external data must be loaded through file I/O (wang2024executable). This makes it difficult to inject pre-existing Python objects (e.g., in-memory DataFrames, trained models) and risks context explosion with complex artifacts (liu2024lost, packer2023memgpt).
@@ 5.3 Context Management & Stateful Architectures
packer2023memgpt introduced MemGPT, an OS-inspired virtual context management system with tiered memory for long-horizon tasks. qiao2023taskweaver proposed TaskWeaver, a code-first framework preserving data structures across turns. However, existing approaches rely on RAG or textual summarization, inherently lossy methods that strip complex runtime objects of structural integrity and executable properties. CaveAgent uses Variable Injection to treat the Python runtime itself as high-fidelity external memory, allowing variables to persist in their native object form without re-tokenization overhead.
5.4 Multi-Agent Coordination
li2023camel proposed CAMEL for role-playing cooperation; qian2023chatdev introduced ChatDev with chat-chain workflows; hong2023metagpt developed MetaGPT encoding SOPs into prompts. These frameworks rely on text-based message passing, introducing serialization bottlenecks when transferring complex state between agents. CaveAgent enables Runtime-Mediated State Flow: agents collaborate by directly injecting and retrieving variables in a shared runtime, shifting coordination from “communication by talking” to “communication by shared state.”
Figure 6 illustrates the evolution of paradigm shifts in agentic tool use that motivates our work.
@@ 6 Acknowledgment
  • We thank Rui Zhou, a professional UI designer at Metasequoia Tech, for his assistance with the figure design in this paper.
  • We thank Qiuyang Mang, a Ph.D student in Computer Science at UC Berkeley, for the discussion about the core design of our framework.
@@ Appendix B What Happens in Semantic Stream
The following sections detail the prompt templates used to instruct the Semantic Stream in CaveAgent. The system prompt is dynamically constructed by combining the Agent Identity, Context Information (functions, variables, types), and Instructions.
@@ Appendix C What Happened in Runtime Stream
While the Semantic Stream governs reasoning and planning, the Runtime Stream serves as the execution engine and persistent memory. This stream operates as a dedicated Python kernel where data manipulation, tool invocation, and state transitions occur. The two streams follow a strict chronological topology, synchronized through interleaved exchange of code instructions and execution feedback.
@@ C.1 Environment Initialization via Injection
The runtime lifecycle begins with Context Injection. Before the reasoning cycle starts, the user (or the system orchestration layer) initializes the runtime environment by injecting native Python objects directly into the global namespace.
@@
  • Variable Injection: Domain-specific data, such as DataFrames, graph structures, or class instances, are instantiated within the runtime stream’s memory.
@@ C.3 Illustrative Case Study
Viewing the runtime stream as a Jupyter notebook with multiple cells, where each cell corresponds to the execution at each time step, helps illustrate how states remain persistent across steps.
@@ D.1 Type Proficiency Cases
The Type Proficiency category evaluates the agent’s ability to perform precise, state-aware manipulation of Python runtime elements. This section tests the agent’s working memory across three structural tiers: Simple Types (primitive types such as lists, dictionaries, and strings), Object Types (custom classes), and Scientific Types (high-dimensional complex data). Proficiency in these domains is a prerequisite for complex reasoning tasks.
@@ D.3 Multi-turn Cases
These test cases evaluate the agent’s capability to process sequential instructions and maintain state precision over long-horizon scenarios. Unlike single-turn tasks where information is self-contained, these scenarios require the agent to maintain persistent memory of the system’s status, as subsequent queries depend on the outcome of previous actions. We categorize these multi-turn benchmarks into two domains: Smart Home Control and Financial Account Management.
@@ D.3.1 Smart Home
As illustrated in Figure 12, the weekend_party case spans a simulated 24-hour cycle. The agent must maintain a coherent environment state, transitioning from a quiet morning to a loud party and finally to a secure night mode, without drifting from the user’s cumulative intent.
@@ D.3.2 Financial Account
The Financial Account benchmark evaluates the agent’s capability to maintain strict numerical integrity and execute state-dependent logic within a banking ledger system. Unlike the relative adjustments in Smart Home, this domain demands exact integer arithmetic, where the agent must process a continuous stream of transactions, including deposits, interest applications, and loan amortizations, without cumulative drift.
This scenario imposes two constraints designed to stress-test the agent’s reasoning stability:
@@ Appendix E Stateful Runtime-Mediated Multi-Agent Coordination
The function-calling paradigm in CaveAgent introduces three key contributions for multi-agent coordination; Figure 1 illustrates an example. In this paper, we focus on qualitative analysis and provide case studies to facilitate understanding, leaving rigorous quantitative evaluation for future work. We introduce the high-level ideas below.
@@ E.1 Meta-Agent Runtime Control
Sub-agents are injected as first-class objects into a meta-agent’s runtime, enabling the meta-agent to programmatically access and manipulate child agent states through generated code. Rather than following predefined communication protocols, the meta-agent dynamically sets variables in sub-agent runtimes, triggers execution, and retrieves results, enabling adaptive pipeline construction, iterative refinement loops, and conditional branching based on intermediate states.
@@ E.2 State-Mediated Communication
Inter-agent data transfer bypasses message passing entirely. Agents communicate through direct runtime variable injection: the meta-agent retrieves objects from one agent’s runtime and injects them into another’s as native Python artifacts (DataFrames, trained models, statistical analyses), preserving type fidelity and method interfaces without serialization loss.
@@ E.3 Shared-Runtime Synchronization
For peer-to-peer coordination, multiple agents can operate on a unified runtime instance, achieving implicit synchronization without explicit messaging. When one agent modifies a shared object, all peers perceive the change immediately through direct reference. New entities injected into the shared runtime become instantly discoverable, enabling collaborative manipulation of a unified world model with low coordination overhead.
How the town simulation demonstrates this capability. When the meta-agent modifies the weather state, all resident agents observe the change through direct attribute access; when a new location and manager are injected, existing agents can immediately query and interact with them.
Together, these patterns transform multi-agent systems from lossy text-based message exchange into typed, verifiable state flow, enabling automated validation of inter-agent handoffs and integration with downstream pipelines.
E.4 AutoML Training Loop
We demonstrate CaveAgent’s hierarchical agent coordination through an AutoML training loop where an orchestrator agent programmatically manages sub-agent runtimes (Figure 14). The orchestrator injects raw data into a feature engineering agent’s runtime via inject(), triggers execution, then retrieves the transformed DataFrame via retrieve() and injects it into a trainer agent’s runtime, all as native Python objects without serialization. After training, the orchestrator extracts evaluation metrics directly from the trainer’s runtime, validates against target requirements via a check_requirements() function, and injects performance feedback back into both sub-agents for the next iteration. Crucially, sub-agents themselves are injected as variables into the orchestrator’s runtime, enabling the orchestrator to dynamically access and manipulate their internal states through generated code. This iterative refinement loop continues until programmatic convergence criteria are met, demonstrating CaveAgent’s unique capability for hierarchical multi-agent coordination with typed, bidirectional state flow and automated convergence verification.
Figure 14: Multi-agent AutoML pipeline: an orchestrator coordinates feature engineering and training sub-agents through runtime variable injection. Native Python objects (DataFrames, trained models) flow losslessly between agents across iterative refinement cycles.
Appendix F Detailed BFCL Benchmark Results
Table F shows the detailed per-run results of BFCL benchmark.
Table 7: Detailed Performance comparison on BFCL Benchmark (3 Runs). Data is presented in score/total format. The Avg. columns indicate the average overall percentage across 3 runs. Bold indicates CaveAgent outperforms Function Calling. Simp. = Simple Function, Mult. = Multiple Function, Para. = Parallel Function, P-M. = Parallel Multiple Function, Ov. = Overall.
Model Run Function Calling CaveAgent
Simp. Mult. Para. P-M. Ov. Avg.(%) Simp. Mult. Para. P-M. Ov. Avg.(%)
Open Source
Closed Source
R3 387/400 190/200 184/200 184/200 945/1000 94.4 385/400 189/200 184/200 186/200 944/1000 94.4 ($\scriptstyle\sim$0.0)
R3 384/400 190/200 188/200 182/200 944/1000 94.3 380/400 194/200 184/200 185/200 943/1000 94.3 ($\scriptstyle\sim$0.0)
Appendix G Features
G.1 Case Analysis in Tau2-bench
To validate the architectural advantages of CaveAgent, we analyzed trajectory differences on the Tau2-bench retail benchmark. CaveAgent achieved a 72.8% success rate (83/114) compared to 62.3% (71/114) for the baseline JSON agent (Kimi K2 backbone), yielding a 10.5% improvement. We conducted a root cause analysis on the 24 tasks where CaveAgent succeeded but the baseline failed.
G.1.1 Failure Taxonomy of the Baseline
@@
G.1.2 Architectural Advantages: Loops and Conditionals
The analysis reveals that CaveAgent’s improvements stem from its ability to generate programming constructs, specifically loops (used in 92% of winning cases) and conditionals (83%), which resolve the semantic gaps inherent in single-step function calling.
@@ G.1.2 Architectural Advantages: Loops and Conditionals
G.2 Smart Home
Figure 16 illustrates the mechanistic advantage of CaveAgent through a toy smart-home example. The architecture separates the Semantic Stream (logic generation) from the Runtime Stream (state storage). This design enables two key capabilities absent in standard JSON agents:
@@ G.2 Smart Home
Figure 16: Demonstration of CaveAgent in a smart-home example: the Semantic Stream interacts with the Runtime Stream by generating code that manipulates stateful objects (variables) in the persistent runtime.
G.3 Geospatial Analysis
To illustrate CaveAgent’s advantages in domain-specific applications, we consider an urban planning scenario in which a user draws two arbitrary regions on an interactive map and queries the differences in population density and urban development between them. The interactive map converts user-drawn regions into GeoJSON polygon objects, variable-length coordinate arrays with high-precision floating-point pairs, which are injected directly into CaveAgent’s persistent runtime as first-class Python variables. Upon receiving the natural language query, the LLM generates a single code block that references these injected geometries to perform zonal statistics against WorldPop raster data and extract land use features from OpenStreetMap via osmnx, chaining multiple domain-specific operations through native variable passing. Under the conventional JSON function calling paradigm, the same task would require at least five sequential LLM turns, querying population and land use statistics separately for each region before synthesizing text-serialized results, while also confronting the challenge of encoding complex polygon geometries as JSON string parameters, which risks truncation and introduces serialization overhead. CaveAgent resolves the entire query in a single turn with lossless data flow: geometries maintain full numerical precision, intermediate results (e.g., GeoDataFrames) persist as manipulable runtime objects, and the LLM synthesizes the final response from deterministic execution output (Figure 17). This case study demonstrates CaveAgent’s suitability for scientific and analytical domains where computation involves complex non-serializable data structures and precision-sensitive results.
Figure 17: Geospatial Analysis: a user draws two regions on an interactive map. CaveAgent injects the GeoJSON polygons as Python variables and resolves the entire spatial query in a single turn with lossless data flow.